# OSS Malicious Package Analysis in the Wild

### v2

@@ Abstract
The open-source software (OSS) ecosystem suffers from various security threats and risks, and malicious packages play a central role in software supply chain (SSC) attacks. Although malware research has a history of over thirty years, less attention has been paid to OSS malware. Its existing research has three limitations: a lack of high-quality datasets, malware diversity, and attack campaign context. In this paper, we first build and curate the largest dataset of 23,425 malicious packages from scattered online sources. We then propose a knowledge graph to represent the OSS malware corpus and conduct malicious package analysis in the wild. Our main findings include (1) it is essential to collect malicious packages from various online sources because there is little data overlap between different sources; (2) despite the sheer volume of SSC attack campaigns, many malicious packages are similar, and unknown/sophisticated attack behaviors have yet to emerge or be detected; (3) OSS malicious package has its distinct life cycle, denoted as {changingreleasedetectionremoval}, and slightly changing the package (different name) is a widespread attack manner; (4) while malicious packages often lack context about how and who released them, security reports disclose the information about corresponding SSC attack campaigns.
I Introduction
@@
Open-source software (OSS) ecosystems provide a large number of reusable packages, libraries, tools, and processes for software project development. Meanwhile, OSS ecosystems suffer from a variety of security threats and risks, called software supply chain (SCC) attacks [1], where malicious code is injected into the development, build, and release of artifacts. Consequently, OSS malware, software packages containing malicious code segments that execute unauthorized/malicious behaviors, plays a central role in SSC attacks. Malicious packages initially deceive developers and users to download and install them and then execute subsequent behaviors, such as implanting backdoors [2], stealing sensitive information [3], and downloading and executing payloads without user permission (e.g., cryptominers [4]). For example, the ‘Fallguys’ [5] is a malicious package that steals sensitive data stored on a computer, extracts the AWS token key, and sends it to a remote server. This software package has been downloaded more than 345 times, posing a significant security threat to users. Therefore, studying the OSS malicious package and building a defense for safeguarding the OSS ecosystem is highly needed.
@@
OSS malware has been studied in recent years, with focus on the detection and analysis of malicious packages [6, 7, 8, 9, 10]. However, the existing malware research has several limitations. First, traditional malware studies [11, 12, 13, 14, 15] provide compiled malware in the binary format rather than OSS malicious packages. Most malicious packages use interpreted code, and there are no binary-level malware samples. Second, the existing OSS malware datasets [16, 17, 18] are limited in size and from a single OSS ecosystem: 1,195 malicious packages in the GitHub repository [16], 2,915 PyPI malicious packages in [17], and only 200 malicious packages in [18]. Third, many malicious packages are unavailable because the registry administrators remove the packages. As per our investigation, the missing rate is close to 64.14%. Fourth, the diversity of malicious packages is not considered in the existing datasets. A large number of malicious packages does not imply malware diversity. For example, we only obtain 25 code groups from the prior PyPI malware dataset (2,915).
@@
To address those limitations, we first collect a large-scale and high-quality dataset of OSS malware packages from different online sources. Our observation is that security professionals detect and disclose malicious packages. Many commercial security enterprises have their own security blogs to report their findings and malware analysis. Those security enterprises regularly scan the newly added software packages in the OSS ecosystems and discover new and unknown malicious packages. Once publicly disclosed, the analysis information about a malicious package is freely available to a third party. Based on this observation, we have collected 23,425 OSS malicious packages covering 10 different OSS ecosystems. To our knowledge, this is the largest OSS malware dataset.
@@
Further, we leverage the knowledge graph to represent the OSS malware dataset, where each node is a malicious package, and each edge represents a package relationship. The conventional malware family is designed to demonstrate the diversity of binary samples, unsuitable for OSS malicious packages. To resolve this issue, we propose four types of edges (Section III-B): duplicated, dependency, similar, and co-existing relationships. Malicious packages are grouped into a subgraph to represent an OSS malware family, indicating the malware diversity. To launch an attack campaign, a threat actor will release a malicious package multiple times. We also use the subgraph to characterize an attack campaign towards OSS ecosystems and provide a quantitative analysis of the attack campaign along with time.
In this paper, we conduct an empirical study of malicious packages in the wild, aiming to answer the following research questions (RQs):
RQ1: What impact does ad-hoc research have on malicious package datasets?
RQ2: How does the knowledge graph perform when applied to represent the diversity of malicious packages?
RQ3: What information of attack campaigns can a malicious packet group reflect?
RQ4: How do the malicious packages evolve in an attack campaign over time?
@@
Findings. Based on our analysis results, there are four major learned lessons. (1) There is a lack of industry-wide and industry-academia collaboration, limiting the data quality of malicious packages. There is little data overlap, and many studies are conducted independently. OSS malware analysis research is still at a fancy stage, and collecting malicious packages from various online sources is imperative. (2) Despite an increasing number of OSS malicious packages, their diversity evolution is slow and relatively stable. This suggests that today’s defense tools work well because malicious packages use old and known attack behaviors. (3) OSS malicious package has its distinct life cycle, denoted as {changingreleasedetectionremoval}. An SSC attack campaign is a repeating process in which attackers change malicious packages to launch several attack attempts. Slightly changing the package (different name) is a widespread attack method; dependency attack methods are rare but have the longest active period. (4) Malcious packages often lack context about how and who released them, but security reports reveal missing SSC attack campaigns.
Roadmap. Section 2 presents our data collection methodology. Section 3 describes the knowledge graph in our study. Section 4 demonstrates the empirical studies of OSS malicious packages. Section 5 surveys the related work, and finally, Section 6 concludes.
II Data Collection Methodology
@@
In this section, we collect malicious packages from the Internet based on the following three observations. (1) Some known/public malicious package datasets are available, and we can merge them into a large dataset. (2) When malicious packages are not available, we use their names/versions to find corresponding packages. (3) Many package mirror registries are not synced with the root registry, where malicious packages may still exist in the mirror registry. We can find malicious packages in the mirror registry via package names/versions. Figure 1 shows the overall of our data collection approach. We relied on four types of online sources to find public malicious packages, including open-source datasets, commercial websites, social networks (SNS), and individual blogs. In the remainder of this section, we present the details of our data collection process.
II-A Open-source Dataset
@@
To collect malicious packages from the research paper, we conducted an in-depth study of relevant literature about the OSS ecosystem and the SCC attack. Typically, we have found four available open-source datasets for malicious packages, including the Maloss dataset [16], Backstabber-Knife dataset [19], Mal-PyPI dataset [17], and DataDog dataset [20]. For those open-source datasets, we directly download malicious packages if they are available.
@@
A practical problem for those open-source datasets is that the package number is inconsistent with the number in the research work. Mal-PyPI Dataset [17] claimed it collected 4,500+ PyPI packages, but we only obtained 2,915 malicious packages. The remaining 1,500+ packages are unavailable, and we cannot obtain their package names and versions. To deal with the practical issue, we record the package name/version from the open-source dataset if malicious packages are unavailable. We use mirror registries to search package names/versions for obtaining their packages detailed in Section II-C.
II-B Web Crawler
To collect malicious packages, we explored security analysis reports and found three typical online sources, including commercial security websites, social networks, and individual blogs. Security researchers or professionals often share malicious package information and publish security reports in online sources. The most significant limitation of online sources is that they don’t provide any package to the public because they treat malicious packages as important assets and do not share the dataset. Fortunately, they provide the package names and versions in the security reports.
@@ II-B Web Crawler
Commercial Websites and Individual Blogs publish security reports about the malicious packages that they detected or found in the OSS ecosystem. Typically, a webpage about a security analysis report contains several parts: (1) when or how to find the malicious package, (2) its malicious behaviors or objectives, (3) the package name and version, and (4) a possible useful indicator of compromised (IOC). For instance, Snyk.io reported the backdoor packages [21] in the RubyGem ecosystem, where its name is ‘bootstrap-sass’ and its version is ‘Version 3.2.0.3’. We investigated the top 1 million websites in the Alax rank and picked up popular websites related to OSS ecosystems and SCC attacks. From these sources, we manually picked those related to OSS ecosystems and added them to the list, including some research blog websites with software security reports. Our source lists are follows: Snyk.io [22], Tianwen [23], GitHub Advisory [24], Phylum [25], Socket [26], and individual blogs [27, 28, 29]. We use a web crawler to collect webpages from those online sources. The input of our web crawler is the website URL, and the output is the HTML pages. Specifically, we used keywords (e.g., ‘malicious’ and ‘malware’) to filter out irrelevant HTML pages. Once we obtain their webpages, we leverage the BeautifulSoup [30] to parse the HTML tags of the webpages and extract report content. We manually inspect the report content and identify the package names and versions. We record the package names and versions in the list.
@@
Social Networks (SNS) are the source for collecting malicious packages. From our observation, the X (Twitter) is the most active platform with the highest update frequency for finding malicious packages. For instance, we followed an X (twitter) account named “@sscblog” [31] and found that it had released 1.7 malicious packages per day.
For the X (Twitter), we use its API to obtain the tweets and parse their HTML tags. Specifically, we used keywords (e.g., ‘malware’ and ‘SSC’) to pick up popular Twitter accounts that disclose malicious package information. Then, we follow the tweet history of those accounts to collect package names. If the Twitter account points to the same organization or individual from commercial websites and individual blogs, we merge their package names/versions. Note that many SNS tweets are the same as the webpage content on commercial websites and individual blogs. If the Twitter account points to an unconvincing organization or individual, we also attempt to collect those packages but do not add them to the dataset.
II-C Exploring Mirror Registry
@@
We use package names/versions as keywords to search a mirror registry. Malicious packages are removed from the package registry but are still available in the package mirror registry. The reason is that the package mirror registries are not synced with the root registry, and there is a time gap between them. If they exist, we download them from the package registry mirrors. If it does not exist in the registry mirrors, we only have the package name and version without its package. Specifically, we use 5 NPM mirrors [32, 33, 34, 35, 36], 12 PyPI mirrors [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and 6 RubyGems mirrors [49, 50, 51, 52, 53, 54] to search for malicious packages.
@@
Table I lists sources and sizes of initial malicious packages. In total, we have collected 23,730 malicious packages, 9,141 of which are available and 14,589 of which are unavailable. The dataset covers 10 OSS ecosystems, including PyPI, NPM, RubyGems, Maven, Cocoapods, SourceForge, Docker, Composer, NuGet, and Rust. One straightforward observation is that most malicious packages come from PyPI and NPM, popular OSS ecosystems where criminals and adversaries prefer to launch attacks.
Category | Data Source | Unavailable Num # | Available Package # |
Academia | Backstabber-Knife [19] | 3,928 | 1,025 |
Maloss [16] | 3 | 1,195 | |
@@
Snyk.io [22] | 1,161 | 384 | |
Tianwen [23] | 1,772 | 1,429 | |
DataDog [20] | 0 | 1,397 | |
Phylum [25] | 6,669 | 642 | |
@@
Individual | SNS/Blogs [27, 28, 29] | 59 | 3 |
Total | 14,422 | 9,003 |
II-D Dataset Validity
To guarantee the reproducibility of our malicious package analysis, we use the requirements by Rossow et al. [55] to ensure the correctness and transparency of our dataset.
Dataset Correct. We assume that online sources are reliable and trustworthy and have a good reputation in the security community. If a legitimate package is falsely reported as a malicious package, it may be added to contain our dataset. Due to the large number of malicious packages, manually inspecting each package via the reserved engineering technique is impractical. We use a simple rule to remove the false positives: if the root register does not remove a package, it is not a malicious package. Further, we release our dataset via the GitHub repository, where we can receive feedback from the community to remove the false positives.
Dataset Diversity. A large number of unique malicious packages does not imply diversity, where many packages are similar. Diversity is an important indicator for measuring the quality of malicious package datasets. If many malicious packages are centered around a few ones, the dataset is not representative, and the aggressive malicious packages dominate the dataset. Prior works [16, 18, 17] didn’t take into account the diversity of the malicious packages. We propose the knowledge graph to address the diversity of malicious packages detailed in Section III-B.
Dataset Transparency. To guarantee the reproducibility of analysis results, we provide the details of the transparency of the malicious package dataset. Due to page limitations, we present only a summary of our dataset in the paper; the detailed dataset is available in the GitHub repository. We build a website to publish all malicious package names (sources) with their signatures (e.g., MD5 hashes) in our paper. We also list all package groups (manual labeling) so the researcher can identify which package to use in the dataset. Our analysis contains all malicious packages from Table I.
Dynamic Changing. One concern is that the analysis results may be changed when new and unknown malicious packages are released or missing packages are supplemented. The release timeline of the malicious packages is shown in Figure 2, covering from 2018 to 2024. We plot the release time of all malicious packages, with the X-axis showing the release time and the Y-axis showing the package number. Our dataset covers an extended period, and the analysis results are stable with time.
@@ III MalGraph: Malicious Package Graph
This section proposes a knowledge graph (MalGraph) to represent the malicious packages and their relationships. Figure 3 shows the malicious package graph from our collected dataset. A graph is denoted as , where the is the set of nodes, and is the set of edges. Each node represents a malicious package, and each edge represents the relationship between two malicious packages. We propose four types of edges to represent the relationship between two packages, including duplicated, dependency, similar, and co-existing relationships. A graph stores interlinked nodes as a knowledge graph, where we use a graph-structured data model to analyze malicious packages and their relationships.
III-A Node & Edge
Each malicious package is stored as a node in the graph. Specifically, there are seven attributions for the node, including the ID, the package name, the package version, the source, hash value, the path, and the ecosystem.
Duplicated Edge indicates two malicious packages belong to the same package. Specifically, we use the package’s signature to identify whether two packages have a duplicate relationship. If the malicious package is unavailable, we use the package name and version to determine the duplicated edge. For instance, ‘brock-loader-1.9.9’ from Maloss [56] and ‘soltalabs-ramda-extra-1.99.99’ from Tianwen [23] have the same Hash value, denoted as the duplicated relationship ‘brock-loader-1.9.9’, ‘soltalabs-ramda-extra-1.99.99’.
Dependency Edge indicates a malicious package relies on another malicious package. We have extracted all dependency libraries from each package’s metadata file, which is stored as the configuration file, e.g., ‘*.json’ for NPM and ‘*.requirement’ for PyPI. Note that malicious packages may claim legitimate packages as their dependency libraries. We remove those dependency libraries from legitimate packages, only considering the dependency between malicious packages. For instance, ‘loglib-modules’ and ‘pygrata-utils’ are two malicious packages that rely on a malicious package called ‘pygrata’ [57], denoted as the dependency relationship ‘loglib-modules’, ‘pygrata’ and ‘pygrata-utils’, ‘pygrata’.
Similar Edge indicates two malicious packages use a similar source code base. We extract each package’s source code and calculate the similarity between two malware codes using the following process. (1) The source code is converted into an abstract syntax tree (AST), which represents the structure of a program or code snippet. (2) Each AST is converted into an embedded vector using the OPENAI pre-training model. The maximum token length of the code is 8,000, and the average size of the embedded vector is 3,072. (3) The K-Means algorithm clusters the embedded vectors into similar groups. The initial number of clusters is set to 3, and we increase the number of clusters until the centroids of newly formed clusters do not change. (4) We build a similar edge between two malicious packages in the same group. For instance, ‘pylibsql’ has identical code to ‘pylibfont’ package, denoted by the similar relationship ‘pylibsql’, ‘pylibfont’.
Node | Edge | Ave. OutDegree | Ave. InDegree | |
DG | 2,475 | 316,122 | 127.72 | 127.72 |
DeG | 28 | 16 | 0.57 | 0.57 |
SG | 6,320 | 5,343,792 | 845.54 | 845.54 |
CG | 2,941 | 575,406 | 195.65 | 195.65 |

### v3

# An Analysis of Malicious Packages in Open-Source Software in the Wild

@@ Abstract
The open-source software (OSS) ecosystem suffers from security threats caused by malware. However, OSS malware research has three limitations: a lack of high-quality datasets, a lack of malware diversity, and a lack of attack campaign contexts. In this paper, we first build the largest dataset of 24,356 malicious packages from online sources, then propose a knowledge graph to represent the OSS malware corpus and conduct malware analysis in the wild. Our main findings include (1) it is essential to collect malicious packages from various online sources because their data overlapping degrees are small; (2) despite the sheer volume of malicious packages, many reuse similar code, leading to a low diversity of malware; (3) only 28 malicious packages were repeatedly hidden via dependency libraries of 1,354 malicious packages, and dependency-hidden malware has a shorter active time; (4) security reports are the only reliable source for disclosing the malware-based context.
Index Terms:
Malicious Packages, Software AnalysisI Introduction
@@
Open-source software (OSS) ecosystems provide a large number of reusable packages, libraries, tools, and processes for software project development. Meanwhile, OSS ecosystems suffer from a variety of security threats and risks, called software supply chain (SSC) attacks [1], where malicious code is injected into the development, build, and release of artifacts. Consequently, OSS malware, software packages containing malicious code segments that execute unauthorized/malicious behaviors, plays a central role in SSC attacks. Malicious packages initially deceive developers and users to download and install them and then execute subsequent behaviors, such as implanting backdoors [2], stealing sensitive information [3], and downloading and executing payloads without user permission (e.g., cryptominers [4]). For example, the ‘Fallguys’ [5] is a malicious package that steals sensitive data stored on a computer, extracts the AWS token key, and sends it to a remote server. This software package has been downloaded more than 345 times, posing a security threat to users. Therefore, studying the OSS malicious package and building a defense for safeguarding the OSS ecosystem is highly needed.
@@
OSS malware has been studied in recent years, with a focus on the detection and analysis of malicious packages [6, 7, 8, 9, 10]. However, the existing malware research has several limitations. First, traditional malware studies [11, 12, 13, 14, 15] provide compiled malware in the binary format rather than OSS malicious packages. Most OSS malicious packages use interpreted code, and there are no binary-level malware samples. Second, the existing OSS malware datasets [16, 17] are limited in size and from a single OSS ecosystem: 1,195 malicious packages in the GitHub repository [16], and 2,915 PyPI malicious packages in [17]. Third, many malicious packages are unavailable because the registry removes the packages. As per our investigation, the missing rate is close to 39.27%. Fourth, the diversity of malicious packages is not considered in the existing datasets. A large number of malicious packages does not imply malware diversity. We only obtain 153 similar code groups from the PyPI malware dataset (Section V-B).
@@
To address those limitations, we first collect a large-scale and high-quality dataset of OSS malware packages from different online sources. Our observation is that security professionals detect and disclose malicious packages. Many commercial security enterprises have their own security blogs to report their findings and malware analysis. Those security enterprises regularly scan the newly added software packages in the OSS ecosystems and discover new and unknown malicious packages. Once publicly disclosed, the analysis information about a malicious package is freely available to a third party. Based on this observation, we have collected 24,356 OSS malicious packages. To our knowledge, this is the largest OSS malware dataset.
@@
Further, we leverage the knowledge graph (MalGraph) to represent the OSS malware dataset, where each node is a malicious package, and each edge represents a package relationship. We propose four types of edges (Section III): duplicated, dependency, similar, and co-existing relationships. Malicious packages are grouped into a subgraph via those relationships. MalGraph uses a subgraph to represent the OSS malware-based context (e.g., an attack campaign, or malware diversity). For example, a threat actor will release a malicious package multiple times to launch an attack campaign, and those packages may be grouped together. We also use the subgraph to characterize an attack campaign toward OSS ecosystems and provide a quantitative analysis of the OSS malware-based context.
@@
Findings. Based on our analysis results, there are four major learned lessons. (1) There is a lack of industry-wide and industry-academia collaboration, limiting the data quality of malicious packages. There is little data overlap, and many studies are conducted independently. OSS malware analysis research is still in the infancy stage, and collecting malicious packages from various online sources is imperative. (2) Despite an increasing number of OSS malicious packages, their code diversity is low. A typical attack campaign is a repeated attempt in which attackers reuse similar code to distribute different malicious packages to the registry. (3) A small number of malicious packages are repeatedly concealed within other malicious packages, with some being hidden over 400 times. Yet, using dependency libraries to hide malicious packages has a short average active period (10.5 days), and the repetitive distribution based on similar code remains active for approximately 45.16 days. (4) Malcious packages often lack context about how and who released them, but security reports reveal missing SSC attack campaigns. The reported attack campaign discloses thousands of suspicious IP addresses and malicious domains.
II OSS Malware Collection
A reliable and high-quality dataset of OSS malware is essential for the security community. In this section, we first highlight the distinguishing characteristics of OSS malware and then detail the methodology for collecting the OSS malware dataset.
II-A OSS Malware
OSS malware differs from conventional malware in three aspects. (1) Distribution Channels: OSS and conventional malware use different channels for distribution. OSS malware primarily spreads through package registries (e.g., PyPI and NPM), while conventional malware is typically delivered through phishing emails, malicious attachments, infected software, or compromised websites. (2) Packaging: OSS malware is packaged as a software component within a registry, whereas conventional malware is usually delivered as compiled binaries. This difference in packaging allows OSS malware to be integrated into software libraries, while conventional malware varies by architecture and compiler, resulting in distinct binary samples for different systems. (3) Social Engineering: OSS malware uses a different social engineering tactic to deceive users into downloading and installing it. It often uses names, versions, and descriptions similar to trusted software packages (e.g., typosquatting and combosquatting), making it difficult for victim users to differentiate between legitimate and malicious software.
In terms of the above different places, the life cycle of a malicious package can be characterized by four distinct phases: development, release, detection, and removal, as illustrated in Figure 1. The first phase starts when an adversary or a hacker develops a malicious package. During this phase, social engineering tactics are employed to disguise the malware and make it appear as a legitimate package. The second phase occurs when the adversary releases the malicious package into the OSS ecosystem. This stage determines the distribution channel through which the malware spreads, with the package being added to a registry and remaining in this form throughout the subsequent phases. In the third phase, the malicious package is detected by a security researcher, a security company, or a package management administrator. If the malicious package is first found by a security researcher or a security company, they will notify the package manager administrator. Finally, the package administrator removes the malicious package from the registry. A typical malware’s life can be summarized as the sequence: {developmentreleasedetectionremoval}.
The distinguishing characteristics of OSS malware lead to two practical challenges in malware collection:
Once package managers take down malware, it becomes inaccessible from package registries such as NPM, PyPI, and RubyGems.
Ethical and business considerations prevent companies and security professionals from sharing malware datasets publicly.
II-B Data Collection Methodology
@@
To address these challenges, we propose a data collection methodology, as illustrated in Figure 2. (1) Some known/public malicious package datasets are available, and we can merge them into a large dataset. (2) When malicious packages are not available, we use their names/versions to find corresponding packages. (3) Many package mirror registries are not properly synced with the root registry. This means that malicious packages may still exist in the mirror registry, even if they were removed from the main registry. We can find malicious packages in the mirror registry via package names/versions.
@@
Open-source Dataset. To collect malicious packages for the study, we conducted an in-depth investigation of relevant literature about the OSS ecosystem and the SCC attack. Typically, we have found four available open-source datasets for malicious packages, including the Maloss dataset [16], Backstabber-Knife dataset [18], Mal-PyPI dataset [17], and DataDog dataset [19]. For those open-source datasets, we directly download malicious packages if they are available.
@@
A practical problem for those open-source datasets is that the package number is inconsistent with the number in the research work. Mal-PyPI Dataset [17] claimed it collected 4,500+ PyPI packages, but we only obtained 2,915 malicious packages. The remaining 1,500+ packages are unavailable, and we cannot obtain their package names and versions. To deal with the practical issue, we record the package name/version from the open-source dataset if malicious packages are unavailable. We use mirror registries to search package names/versions to obtain their packages.
@@
Commercial Websites and Individual Blogs. Security researchers or professionals often share malicious package information and publish security reports on webpages. Typically, a webpage about a security analysis report contains several parts: (1) when or how to find the malicious package, (2) its malicious behaviors or objectives, (3) the package name and version, and (4) a possible useful indicator of compromised (IOC). For instance, Snyk.io reported the backdoor packages [20] in the RubyGem ecosystem, where its name is ‘bootstrap-sass’ and its version is ‘3.2.0.3’. We investigated an in-depth investigation of popular websites related to OSS ecosystems and SCC attacks. From these websites, we manually picked those related to OSS ecosystems and added them to the list, including Snyk.io [21], Tianwen [22], GitHub Advisory [23], Phylum [24], Socket [25], and individual blogs [26, 27, 28].
@@
Social Networks (SNS) are the source for collecting malicious packages. From our observation, X (Twitter) is the most active platform with the highest update frequency for finding malicious packages. However, the most significant limitation is that online sources (commercial security websites, social networks, and individual blogs) don’t provide any package to the public because they treat malicious packages as important assets and do not share the dataset.
@@
To deal with the limitation, we use package names/versions as keywords to search multiple mirror registries. Malicious packages are removed from the package registry but are still available in package mirror registries. The reason is that the package mirror registries are not synced with the root registry, and there is a time gap between them. If they exist, we download them from the package registry mirrors. If they do not exist in the registry mirrors, we only have the package name and version without its package. Specifically, we use 5 NPM mirrors [29, 30, 31, 32, 33], 12 PyPI mirrors [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], and 6 RubyGems mirrors [46, 47, 48, 49, 50, 51] to search for malicious packages.
@@
Table I lists sources and sizes of initial malicious packages. In total, we have collected 24,356 malicious packages, 13,932 of which are available and 10,424 of which are unavailable. “Unavailable” indicates we only have package names/versions, while “Available” indicates we have the malware package. The dataset covers 10 OSS ecosystems, including PyPI, NPM, RubyGems, Maven, Cocoapods, SourceForge, Docker, Composer, NuGet, and Rust. One straightforward observation is that most malicious packages come from PyPI and NPM, popular OSS ecosystems where criminals and adversaries prefer to launch attacks. In addition, we find that the overlapping degree between SNS and commercial websites is high. The reason is that many SNS accounts point to the same organizations of commercial websites and individual blogs. Note that those SNS tweets are the same as the webpage content on commercial websites and individual blogs.
Category | Data Source | Unavailable | Available |
Academia | Backstabber-Knife [18] | 0 | 5,937 |
Maloss [16] | 0 | 1,223 | |
@@
Snyk.io [21] | 1,160 | 380 | |
Tianwen [22] | 1,769 | 1,382 | |
DataDog [19] | 0 | 1,387 | |
Phylum [24] | 6,606 | 692 | |
@@
Blogs [26, 27, 28] | 59 | 3 | |
Total | 10,424 | 13,932 |
@@ III MalGraph: Malicious Package Graph
This section proposes a knowledge graph (MalGraph) to represent the malicious packages and their relationships. Figure 3 shows a malicious package graph from our collected dataset. A graph is denoted as , where the is the set of nodes, and is the set of edges. Each malicious package is stored as a node in the graph. Specifically, there are seven attributions for the node, including the ID, the package name, the package version, the source, the hash value, and the ecosystem. We propose four types of edges to represent the relationship between two packages, including duplicated, dependency, similar, and co-existing relationships. Specifically, we use Neo4j [52] to implement the knowledge graph. A graph stores interlinked nodes as a knowledge graph, where we use a graph-structured data model to analyze malicious packages and their relationships.
III-A Duplicated Edge
Duplicated edge indicates two malicious packages are identical packages. We determine whether two malicious packages belong to the same package via the following process. If two malicious packages from different sources have identical names and versions, they are classified as duplicate packages. Next, if malicious packages are available, their hash (e.g., MD5) is calculated to further verify the duplicate relationship. Specifically, we leverage the Hashlib [53] library to calculate the SHA256 of the malware code. For instance, ‘acookie-1.0.0’ from Maloss [54] and ‘acookie-1.0.0’ from Backstabber-Knife [18] have the same name, version and hash value, denoted as the duplicate relationship ‘acookie-1.0.0’, ‘acookie-1.0.0’. In short, a duplicated edge characterizes the same malicious package from different sources, indicating the overlap of malware collection.
III-B Similar Edge
Similar Edge indicates two malicious packages that share a similar source code base but have different signatures. The rationale behind the similar edge is that malware distributors reuse the same malware code to execute a series of repetitive attempts targeting OSS ecosystems. However, once a prior malware package is taken down, the same name cannot be reused, leading to different signatures. As a result, malware distributors must use a different package name for the subsequent attack. Over time, the same distributor may release multiple malicious packages, often based on an identical code base. For example, as shown in Figure 4, a total of 15 malware packages were released between August 9 and 19. Note that we cannot determine whether similar packages are the result of copycat behavior or originate from the same organization or author. Particularly, a similar edge represents a repeated attempt derived from the same malware code base.
We determine a similar relationship between two malicious packages via the following process. (1) Unpacking a malware package into a folder and finding all source code files (.js, .py, .rb). (2) Sorting all files in the same order (file names) and merging source code into a whole file. (3) Splitting the whole file into multiple code snippets, where each snippet has a fixed token length. The embedding model (512 tokes in CodeBERT-base [56]) converts each snippet into a vector. We concatenate all vectors into a large vector to represent the whole package code. (4) The K-Means algorithm clusters malware packages into similar groups based on the large vectors. The minimum similarity threshold for the K-means algorithm is set to 0.7, and each subgraph must contain at least two points. After clustering, the silhouette score is calculated using the Scikit-learn library [57]. If the silhouette score of a cluster is less than 0.3, it is removed, indicating poor clustering performance. (5) We build a similar edge between two malicious packages in the same group. The average intra-group similarity of all subgraphs in the dataset reaches 99.9%.
III-C Dependency Edge
Dependency Edge indicates a malicious package relies on another malicious package. Malware distributors release multiple malicious packages, some of which, pretending to be innocent, use a dependency malicious library to hide their malicious behavior. The benefit is that the malicious package can be shielded by an innocent package. Figure 5 plots the dependent-hidden attack. The malware developers first create a package with malicious codes to the registry repository and then create a second one. The second package, depending on the first one, seems like a legitimate package without malicious behaviors. The malware distributor distributes two packages to OSS registries. The users may download and install the second package from the OSS registry. In the installation stage, the first package is automatically downloaded to the project, and the attack is triggered. For instance, ‘loglib-modules’ and ‘pygrata-utils’ are two malicious packages that rely on a malicious package called ‘pygrata’ [58], denoted as the dependency relationship ‘loglib-modules’, ‘pygrata’ and ‘pygrata-utils’, ‘pygrata’.
We determine a dependency relationship between two malicious packages via the following process. (1) We first extract all package names and versions in the malware dataset. (2) For the metadata, we use dependency libraries stored in the configuration file, e.g., ‘*.json’ for NPM and ‘requirement.txt’ for PyPI. Note that malicious packages may claim legitimate packages as their dependency libraries. If a dependency library is matched, we put them as the same dependent-hidden attack. (3) For the source code, we use all malicious package names to find the matched position in the source file. If matched, we extract the code segment within a 100-character window. We use regular expressions (Table II) to determine whether the code segment contains a dependency library from other malicious packages. (4) We provide manual inspection over those dependency libraries to remove false positives, e.g., code comments. Note that our matching approach belongs to the exact string-matching algorithm.
Pattern |
[…]
require([”\’][^”\’]+(?:/[^”\’]+)*/[”\’]) |
III-D Co-existing Edge